Having some trouble, please Explain your answer thanks.

On January 1, 2021, the Taylor Company adopted the dollar-value LIFO method. The inventory value for its one inventory pool on this date was $310,000. Inventory data for 2021 through 2023 are as follows:

Date Ending Inventory
at Year-End Costs
Cost Index
12/31/2021 $ 339,900 1.03
12/31/2022 382,950 1.11
12/31/2023 398,650 1.19


Required:
Calculate Taylor's ending inventory for 2021, 2022, and 2023.

Inventory Layers Converted to Base Year Cost Inventory Layers Converted to Cost Inventory DVL Cost
Date Inventory at Year-End Cost Year-End Cost Index Inventory Layers at Base Year Cost Inventory Layers at Base Year Cost Year-End Cost Index = Inventory Layers Converted to Cost
01/01/2021 = Base = $0
12/31/2021 = Base =
2021 = $0
12/31/2022 = Base =
2021 =
2022 = $0
12/31/2023 = Base =
2021 =
2022 =
2023 = $0
Answer #1

Answer-

Inventory layers converted to base year cost Inventory layers converted to Cost Inventory DVL Cost
Date Inventory at year end cost / Year end cost index = Inventory layers at base year cost Inventory layers at base year cost * Year end cost index = Inventory layers converted to Cash
01/01/2021 310000 / 1 = 310000 Base 310000 * 1 = 310000
12/31/2021 339,900 / 1.03 = 330000 Base 310000 * 1 = 310000
2021 20000 * 1.03 = 20600 330600
12/31/2022 382,950 / 1.11 = 345000 Base 310000 * 1 = 310000
2021 20000 * 1.03 = 20600
2022 15000 * 1.11 = 16,650 347,250
12/31/2023 398,650 / 1.19 = 335000 Base 310000 * 1 = 310000
2021 20000 * 1.03 = 20600
2022 5000 * 1.19 = 5950
2023 0 * 0 = 0 336,550